Please help me solve this simultaneous Equation : 2x-y = 5
3x- 3y-3/5 =4

Found 2 solutions by rothauserc, Sam2001:
Answer by rothauserc(4718)

1) 2x - y = 5
2) 3x - 3y - (3/5) = 4
***********************************
simplify the second equation
3) 3x - 3y = 23/5
multiply first equation by 3
4) 6x - 3y = 75/5
subtract equation 3) from equation 4)
3x = 52/5
x = 52/15
substitute for x in equation 1)
2*(52/15) - y = 5
104/15 -y = 75/15
y = 29/15
************************************
x = 52/15 and y = 29/15
***********************************
check the answer by substituting for x and y in equation 2)
3(52/15) -3(29/15) = 69/15
156/15 - 87/15 = 69/15
69/15 = 69/15
our answer checks

Answer by Sam2001(1)

Multiply 2x-y = 5 by -3
so, (2x-y = 5)-3
-6x+3y = -15
3x-3y = 23/5
Then, add the two equations
Resulting to -3x = -52/5
Then divide by -3
so, x= 52/15
Substitute x to any equations
2(52/15)-y =5
-y = -29/15
Multiply by -1 so that y becomes positive.
( -y = -29/15 )
Resulting to:
y = 29/15
so answer will be x = 52/15 and y = 29/15